2X-1/X+2 = 4/5
THATS SUPPOSED TO BE 2X-1 OVER X+2 = 4 OVER 5

Answer by tutorcecilia(2152)

2X-1/X+2 = 4/5 [Cross-multiply]
(5)(2X-1)=(4)(X+2)
10x-5=4x+8
10x-4x-5=4x-4x+8
6x-5=8
6x-5+5=8+5
6x=13
x=13/6
.
Check by plugging (x=13/6) back into the original equation.